OpenStudy (anonymous):

[Integrate {sin(sqr. x)/sqr.] x dx Upper limit pi^2 Lower limit 0+

OpenStudy (agreene):

\[\int\limits_{0}^{\pi^2}\frac{\sin(\sqrt{x})}{\sqrt{x}}dx\] is that the question?

OpenStudy (anonymous):

yes, only the 0 has a + sign on top

OpenStudy (agreene):

Yeah, I dont think that really matters...lol So, start by using substitution: Let u= sqrt (x) du = 1/(2 sqrt (x)) and you will see: \[2\int \sin(u) du\] \[-2\cos(u)\] \[-2\cos(\sqrt x)\] evaluate from your limits.

OpenStudy (agreene):

should be 4.